A 0.222 g sample of carbon dioxide, CO2, has a volume of 557 mLand a pressure of 491 mmHg

What is the temperature of the gas in kelvins?


1
Expert's answer
2021-10-26T05:37:20-0400

PV = nRT

Molar mass of CO2= 44.01

0.222/44.01= 0.005mol

0.00131579×491= 0.646atm

557×0.646= 0.005×8.314 J/mol·K×T

359.85=0.04157T

T= 359.85/0.0415

= 8671.08K